# The Galactic population of canonical pulsars II

### v2

@@ Appendix B Results for a constant magnetic field
Hence, this simulation was run with 1 million pulsar simulated, without a death line and without a decaying of the magnetic field. The KS test gave p-values of $1.1\times 10^{-14}$ and $1.5\times 10^{-76}$, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$. The parameters used are in Table 5. It gives results less close to the observations than with a decaying magnetic field. As can be seen in Fig. 25, too many pulsars from the simulation appear on the top right of the diagram and almost none at the bottom. In Table 11, we note a total of pulsars which is smaller from what we found for all of our previous results in Sect. 3. We could expect to reproduce equally well the population of observed pulsars with a constant magnetic field, however compared to the study of Johnston & Karastergiou (2017) where they used this assumption for the magnetic field, we do not evolve the braking index. In our simulations, we have $\dot{\alpha}<0$, $\dot{B}<0$ and $\dot{n}=0$ whereas in their simulation $\dot{\alpha}<0$, $\dot{B}=0$ and $\dot{n}\neq 0$. Having either the braking index or the magnetic field evolve, allows to reproduce well the population of pulsars.

### v3

@@ Appendix B Results for a constant magnetic field
Hence, this simulation was run with 30 million pulsar simulated (meaning the oldest pulsar simulated is $1.23\times 10^{9}$ yr), with a death line and without a decaying magnetic field. The KS test gave p-values of $1.03\times 10^{-22}$ and $3.38\times 10^{-51}$ for $P$ and $\dot{P}$ respectively, therefore the null hypothesis is rejected for both $P$ and $\dot{P}$. Here we simulate more pulsars than in the main part of the paper, since if we simulate with the same number of pulsars, it seems the whole $P-\dot{P}$ diagram is not covered. Therefore if a constant magnetic field was the best way to reproduce the canonical population, it means we would detect much older pulsars than with the magnetic field decay prescription. The parameters used are in Table 5. It gives results less close to the observations than with a decaying magnetic field. As can be seen in Fig. 26, too many pulsars from the simulation appear on the top right of the diagram and almost none at the bottom. In Table 11, we note a total of pulsars which is smaller from what we found for all of our previous results in Sect. 3. We could expect to reproduce equally well the population of observed pulsars with a constant magnetic field, however compared to the study of Johnston & Karastergiou (2017) where they used this assumption for the magnetic field, we do not evolve the braking index. In our simulations, we have $\dot{\alpha}<0$, $\dot{B}<0$ and $\dot{n}=0$ whereas in their simulation $\dot{\alpha}<0$, $\dot{B}=0$ and $\dot{n}\neq 0$. Having either the braking index or the magnetic field evolve, allows to reproduce well the population of pulsars.